evaluate the following integral
∫ cot x / log sin x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{cotx}{log(sinx)}dx
log(sinx) = t
\frac{1}{sinx}cosxdx = dt
cotxdx = dt
I = \int \frac{1}{t}dt
I = log(t) + c
I = log(logsinx) + c